Dividends Per Share

Imaging Inc., a developer of radiology equipment, has stock outstanding as follows: 12,000 shares of cumulative preferred 4% stock, $120 par, and 40,000 shares of $5 par common. During its first four years of operations, the following amounts were distributed as dividends: first year, $38,640; second year, $86,560; third year, $106,640; fourth year, $130,000.

Compute the dividends per share on each class of stock for each of the four years. Round all answers to two decimal places. If no dividends are paid in a given year, enter "0".

1st Year 2nd Year 3rd Year 4th Year
Preferred stock (dividend per share) $ $ $ $
Common stock (dividend per share) $ $ $ $

Selected transactions completed by ATV Discount Corporation during the current fiscal year are as follows:

Jan. 5. Split the common stock 3 for 1 and reduced the par from $75 to $25 per share. After the split, there were 1,150,500 common shares outstanding.
Mar. 10. Purchased 41,400 shares of the corporation’s own common stock at $28, recording the stock at cost.
Apr. 30. Declared semiannual dividends of $0.80 on 74,100 shares of preferred stock and $0.12 on the common stock to stockholders of record on May 15, payable on June 15.
June 15. Paid the cash dividends.
Aug. 20. Sold 29,000 shares of treasury stock at $33, receiving cash.
Oct. 15 Declared semiannual dividends of $0.80 on the preferred stock and $0.12 on the common stock (before the stock dividend). In addition, a 2% common stock dividend was declared on the common stock outstanding. The fair market value of the common stock is estimated at $36. The dividend date of record is November 15 payable on December 19.
Dec. 19. Paid the cash dividends and issued the certificates for the common stock dividend.

Journalize the transactions. If no entry is required, simply skip to the next transaction. Refer to the Chart of Accounts for exact wording of account titles.

CHART OF ACCOUNTS
ATV Discount Corporation
General Ledger
ASSETS
110 Cash
120 Accounts Receivable
131 Notes Receivable
132 Interest Receivable
141 Merchandise Inventory
145 Office Supplies
151 Prepaid Insurance
181 Land
193 Equipment
194 Accumulated Depreciation-Equipment
LIABILITIES
210 Accounts Payable
221 Notes Payable
226 Interest Payable
231 Cash Dividends Payable
241 Salaries Payable
261 Mortgage Note Payable
EQUITY
311 Common Stock
312 Paid-In Capital in Excess of Par-Common Stock
315 Treasury Stock
321 Preferred Stock
322 Paid-In Capital in Excess of Par-Preferred Stock
331 Paid-In Capital from Sale of Treasury Stock
340 Retained Earnings
350 Stock Dividends Distributable
351 Cash Dividends
352 Stock Dividends
390 Income Summary
REVENUE
410 Sales
610 Interest Revenue
EXPENSES
510 Cost of Merchandise Sold
515 Credit Card Expense
520 Salaries Expense
531 Advertising Expense
532 Delivery Expense
533 Selling Expenses
534 Rent Expense
535 Insurance Expense
536 Office Supplies Expense
537 Organizational Expenses
562 Depreciation Expense-Equipment
590 Miscellaneous Expense
710 Interest Expense

Journalize the transactions. If no entry is required, simply skip to the next transaction. Refer to the Chart of Accounts for exact wording of account titles.

How does grading work?

PAGE 10

JOURNAL

ACCOUNTING EQUATION

Score: 211/225

DATE DESCRIPTION POST. REF. DEBIT CREDIT ASSETS LIABILITIES EQUITY

1

2

3

4

5

6

7

8

9

10

11

12

13

14

15

16

17

18
Answer #1

Calculate dividend per share :

Preferred Dividend = 12000*120*4% = $57600

Compute the dividends per share on each class of stock for each of the four years. Round all answers to two decimal places. If no dividends are paid in a given year, enter "0".

1st year 2nd year 3rd year 4th year
Preferred Stock (dividend per share) 38640/12000 = 3.22 76560/12000 = 6.38 57600/12000 = 4.80 57600/12000 = 4.80
Common stock (dividend per share) 0 10000/40000 = 0.25 49040/40000 = 1.23 72400/40000 = 1.81